The distance(d) is directly proportional to time (t) and d = 15 km after 2 1/2 hrs, then after 3 hrs more d is ?
OH GOODNESS PLEASE HELP

Answer by Alan3354(21583)

The distance(d) is directly proportional to time (t) and d = 15 km after 2 1/2 hrs, then after 3 hrs more d is ?
--------------------
In other words, the longer you drive (or run, or move by any means) the farther you go.
distance = rate*time
15 = r*2.5
r = 15/2.5
r = 6 km/hr
d2 = 6*3 (rate*time)
d2 = 18 km